OpenStudy (brittany16):

Solve for the variable. d - (-6) = 15

OpenStudy (anonymous):

d=9

jimthompson5910 (jim_thompson5910):

d - (-6) = 15 d + 6 = 15 d + 6-6 = 15-6 d = 9

OpenStudy (anonymous):

d=9